OpenStudy (anonymous):

Solve by factoring x^2+2x-48=0

OpenStudy (anonymous):

x^2+2x+1=48+1 square root(x+1)^2= square root (49) x+1=7 x=6

OpenStudy (anonymous):

\[x^2 + 2 x - 48 = (x - 6) (x + 8) \]x=6 and x=-8